OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x2 + 2x + 5 = 0.

OpenStudy (anonymous):

no real roots one real root two distinct real roots three distinct real roots

OpenStudy (anonymous):

The discriminant of the equation : \[ax^2+bx+c=0\] is : \[\Delta= b^2-4\times a\times c\] And there is 3 cases : \[1.~~ \Delta >0\] In this case there are two different solutions \[ 2.~~\Delta=0\] In this case there is one sol solution ! \[3.~~\Delta<0\] in this case there are no real solution !

OpenStudy (anonymous):

so which is the answer

OpenStudy (anonymous):

Use the rule above to find the answer !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Did you find the value of : \[\Delta\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is it ?